magine a giant dry-cleaner’s bag full of air at a temperature of −35°C floating like a balloon with a string hanging from it 10 km above the ground. Estimate what its temperature would be if you were able to yank it suddenly back to Earth’s surface.
Answer #1

The combined gas law states that
P V / T = a constant
If the volume constant then P / T = a constant
Where P and T are pressure and the temperature of the gas at a time
The pressure at 10 km above the sea level is P1  = 265 milli bar = 26.5 k Pa
The pressure at the sea level is P2 = 1000 milli bar = 100 k Pa
The temperature of the air at 10 km is T1   = 238 K
We need to find the temperature at the sea level T2
Using the above law
P1 / T1   = P2 / T2
T2  = (P2 / P1) T1
T2 = (100 k Pa / 26.5 ) x 238
T2  = 898.11 K = 625.110C